OpenStudy (anonymous):

Can someone please show me the steps for putting this problem into 1 fraction? (1/x)+(3/y)-(x/xy)

OpenStudy (anonymous):

Start with a common denominator for each fraction

OpenStudy (anonymous):

Create common denominators by selectively multiplying each term by a factor of 1. \[\frac 1x \color{red}{\cdot \frac{y}{y}}+ \frac 3y\color{red}{\cdot \frac{x}{x}} - \frac x{xy}\]\[\frac y{xy} + \frac{3x}{xy} - \frac x{xy}\]There's still one more step you can do to simplify from here.

OpenStudy (anonymous):

@yakeyglee is the correct final answer (2x+y)/(xy)

OpenStudy (anonymous):

Yep, that's right. Good job!